OpenStudy (anonymous):

Explain, in complete sentences, the pattern of exponents on the imaginary number i. What is the usefulness of knowing this pattern when performing operations on complex numbers?

OpenStudy (amistre64):

the pattern is modular, or clock, arithmetic. there are only 4 states that it can take regardless of what the exponent is

OpenStudy (amistre64):

i^0 = 1 i^1 = i i^2 = -1 i^3 = -i

OpenStudy (amistre64):

divide the exponent by 4; and see what your remander is to determine its equivalent value

OpenStudy (anonymous):

and let me add that by "remainder" amistre means whole number remainder. like when you divide 103 by 4 the remainder is 3 i have noticed some confusion on this point